### 15 Dec 2023 v1

@@ 2 Related Work
Generating high-fidelity queries is an important aspect of interactive user simulations [9, 7, 10, 8]. Earlier works propose different query generation approaches that either rely on principled rules or language models. Doc2Query is a method that produces a set of questions that a given document may answer. The first presented use of these questions was to enrich documents during indexing [22]. This enrichment significantly increased the retrieval effectiveness for term-based ranking methods like BM25. [3] have shown that this method is also suitable for varying queries in a simulation environment. Recently, instruction-tuned LLMs were used to generate queries based on a given prompt providing context information about the information need [1]. This approach has not been used for simulating interactive IR sessions, which is a contribution of this work.
@@ 3.1 Retrieval Models
Table 1: Retrieval effectiveness of BM25 (sparse) and MonoT5 (dense).
@@
The retrieval model is a key element in our simulations. In the experiments, we evaluate the benefits of using a more effective model than a baseline ranking. To this end, we compare BM25 [24] against MonoT5 [21] that is known to outperform the former method in a zero-shot setting. To better understand the quality of search results shown to the simulated users, we evaluate the retrieval effectiveness in a preliminary ad hoc retrieval experiment.
Table 1 compares BM25 to MonoT5 at different cutoffs, i.e., the cutoff level determines how many documents are retrieved by the first-stage ranker and will be re-ranked by MonoT5. In principle, the re-ranker can rely on a larger set of potentially relevant documents and place them at higher ranks. However, an increasing cutoff level does not substantially improve P@10 and nDCG@10, which can be explained by unjudged documents placed at higher ranks, as also shown by the increasing Bpref scores of MonoT5.
@@ 3.2 Click and Stop Decisions
To be in line with earlier work, we compare four different types of click behavior in our experiments. All of the methods are based on probabilistic decisions biased towards the level of relevance and are implemented as shown in Table 2. The perfect and almost random browsing behaviors serve as sky- and baselines, respectively. The perfect user always clicks on relevant results and never clicks non-relevant or unjudged documents, i.e., the user scents the relevant information in a strongly idealistic way. In contrast, the almost random behavior mainly neglects the relevance labels. In between, there are the informational and navigational models with different degrees of exploratory behavior. We model the stop decisions in two different ways.
@@ 3.3.1 Prompting query reformulations
Given the information need described by the topic file, we construct a topic-specific prompt for the LLM. More specifically, we let the LLM generate outputs based on the prompt Please generate one-hundred keyword queries about <title>. <description> <narrative>, whereas <title>, <description>, and <narrative> are taken from the respective fields of the particular topic. To better understand the effects of different prompting strategies, we consider the first generation method GPT that omits the additional topic fields <description> and <narrative> in the prompt to have context-free query variants. The second generation method GPT+ uses the entire prompt. We implement two additional methods to better understand the benefits of having different query variants with semantics. GPT* uses the topic’s title as the seed query for each query (re)formulation and expands it with a single term out of the vocabulary generated for the topic by the GPT+ strategy. Like GPT*, GPT** is also a rule-based variant based on the context-free vocabulary of GPT. For each topic in the test collections (cf. 3.5), we adapt the prompting strategy to query OpenAI’s API and parse the outputs of GPT-3.5 (more specifically, gpt-3.5-turbo-0301222https://platform.openai.com/docs/model-index-for-researchers). In total, we generated 400 queries for all 50 topics in two test collections, which resulted in a total of 40,000 generated queries that stem from the outputs of OpenAI’s LLM.
@@ 3.5 Implementation Details and Datasets
We implement the experimental setup with a rich set of state-of-the-art software libraries, including Pyterrier, ir_datasets, SIMIIR v2.0, and HuggingFace’s transformers. We ground the experiments on two TREC newswire test collections, i.e., the New York Times Annotated Corpus [^3] https://catalog.ldc.upenn.edu/LDC2008T19 used as part of TREC Common Core 2017 (Core17) [2], and the TREC Washington Post Corpus [^4] https://trec.nist.gov/data/wapost/ used as part of TREC Common Core 2018 (Core18) [29]. Here it is important to note that Core18 includes $\sim$26k ($\sim$4k positive) relevance labels, while Core19 includes $\sim$30k ($\sim$9k positive) labels. All of the experiments are run on a Dell workstation with an Intel i9-12900K CPU, 64GB of RAM, and an NVIDIA GeForce RTX 3070 GPU on Ubuntu 22.04 LTS.
@@ 5 Discussion
The results in Section 4.1 show that although the dense retrieval model gives slightly better results, these differences stay the same throughout the simulation. Minor differences in trends align with other studies [12, 28], and in our perspective, user behavior leads to a much greater variance in effectiveness. In particular, taking into account the context within the search session — modeled by dynamic click behavior, background knowledge, and the integration of feedback — substantially impacts information gain. From our point of view, it is crucial for future simulation experiments to integrate the user context instead of only investigating cost-intensive retrieval models.
@@
For both test collections, the trends were similar in all evaluations, even though the absolute information gain differed. We suspect this is related to the substantial difference in the number of documents rated as relevant (cf. 3.5).
In future evaluations, the costs for specific user actions used for the effort-based measure should be specified and justified. Still, the order and trends of the effort-based evaluations resemble the sDCG evaluations closely. Furthermore, the effort-based measure can reveal effects that would remain hidden if only the number of queries were considered. For example, Figure 3 shows that the difference between 50tnr and 20rpp is more pronounced for the effort-based measure since the larger number of examined documents comes with high costs for the user that sDCG does not consider.
@@
Regarding sRBP, sDCG also mostly supports the produced orders of the different configurations. However, the information gain stagnates more quickly, and a difference in the orders can be seen in Figure 3, so dynamic click behavior is penalized by sRBP. Both findings can be explained by the substantially stronger discounting of late results by sRBP. Interestingly, dynamic stop behavior is notably rewarded by the other measures. As Figure 4 indicates, the larger number of examined documents can likely further be attributed to this opposing evaluation behavior. Determining not only realistic durations for user actions but also estimating suitable parameters for sRBP is vital for future work to enable more accurate evaluations and could be done by analyzing real user behavior.
@@
The use of probabilistic query generation methods shows promising results. Not only do they dominate rule-based approaches, but they are also more capable of integrating background knowledge of the topic. Integrating feedback into probabilistic approaches would be an exciting direction for future work, as our rule-based approach has clearly shown that feedback is beneficial.

### 19 Jan 2024 v2

@@ 2 Related Work
Generating high-fidelity queries is an important aspect of interactive user simulations [9, 6, 11, 7]. Earlier works propose different query generation approaches that either rely on principled rules or language models. In recent times, there has been a notable focus on leveraging LLMs for query generation [21, 35, 36]. Doc2Query is a method that produces a set of questions that a given document may answer. The first presented use of these questions was to enrich documents during indexing [27]. This enrichment significantly increased the retrieval effectiveness for term-based ranking methods like BM25. [10] have shown that this method is also suitable for varying queries in a simulation environment. Recently, instruction-tuned LLMs were used to generate queries based on a given prompt providing context information about the information need [1]. This approach has not been used for simulating interactive IR sessions, which is a contribution of this work.
@@ 3.1 Retrieval Models
Table 1: Adhoc retrieval effectiveness of BM25 and BM25+MonoT5 evaluated with the New York Times Annotated Corpus (Core17) and the TREC Washington Post Corpus (Core18) that are part of the experimental setup (cf. 3.5). The cutoff values denote the number of documents reranked by MonoT5.
@@
The retrieval model is a key element in our simulations. In the experiments, we evaluate the benefits of using a more effective model than a baseline ranking. To this end, we compare BM25 [29] against MonoT5 [26] that is known to outperform the former method in a zero-shot setting. To better understand the quality of search results shown to the simulated users, we evaluate the retrieval effectiveness in a preliminary test collection-based adhoc retrieval experiment.
Table 1 compares BM25 to MonoT5 at different cutoffs, i.e., the cutoff level determines how many documents are retrieved by the first-stage ranker and will be re-ranked by MonoT5. It is important to emphasize that MonoT5 works according to the retrieve and rerank paradigm, which means that in our case, a set of documents is retrieved in the first phase with a sparse model (BM25 in our case) and reranked in a second phase by a dense model. In principle, the re-ranker can rely on a larger set of potentially relevant documents and place them at higher ranks. However, an increasing cutoff level does not substantially improve P@10 and nDCG@10, which can be explained by unjudged documents placed at higher ranks, as also shown by the increasing Bpref scores of MonoT5.
@@ 3.2 Click and Stop Decisions
To be in line with earlier work [14], we compare four different types of click behavior in our experiments. All of the methods are based on probabilistic decisions biased towards the level of relevance and are implemented as shown in Table 2. We note that other ways exist to simulate clicks [8], when there are historical interaction logs, which unfortunately were not available for our experimental setup. The perfect and almost random browsing behaviors serve as sky- and baselines, respectively. The perfect user always clicks on relevant results and never clicks non-relevant or unjudged documents, i.e., the user scents the relevant information in a strongly idealistic way. In contrast, the almost random behavior mainly neglects the relevance labels. In between, there are the informational and navigational models with different degrees of exploratory behavior. We model the stop decisions in two different ways.
@@ 3.3.1 Prompting query reformulations
Given the information need described by the topic file, we construct a topic-specific prompt for the LLM. More specifically, we let the LLM generate outputs based on the prompt template given below.
Prompt template for the query strategies “GPT” and “GPT+”
Please generate one-hundred keyword queries about <title>. <description> <narrative>
<title>, <description>, and <narrative> are taken from the respective fields of the particular topic. To better understand the effects of different prompting strategies, we consider the first generation method GPT ( ) that omits the additional topic fields <description> and <narrative> in the prompt to have context-free query variants. The second method GPT+ (  ) makes use of the entire prompt to generate query reformulations.
We implement two additional methods to better understand the benefits of having different query variants with semantics. GPT* uses the topic’s title as the seed query for each query (re)formulation and expands it with a single term out of the vocabulary generated for the topic by the GPT+ (  ) strategy. Similarly, GPT** is also a rule-based variant based on the context-free vocabulary of GPT ( ). For each topic in the test collections (cf. 3.5), we adapt the prompting strategy to query OpenAI’s API and parse the outputs of GPT-3.5 (more specifically, gpt-3.5-turbo-0301222https://platform.openai.com/docs/model-index-for-researchers). In total, we generated 400 queries for all 50 topics in two test collections, which resulted in a total of 40,000 generated queries that stem from the outputs of OpenAI’s LLM.
@@ 3.5 Implementation Details and Datasets
We implement the experimental setup with a rich set of state-of-the-art software libraries, including Pyterrier [20], ir_datasets [19], SIMIIR v2.0 [38], and HuggingFace’s transformers [37]. We ground the experiments on two TREC newswire test collections, i.e., the New York Times Annotated Corpus [^3] https://catalog.ldc.upenn.edu/LDC2008T19 used as part of TREC Common Core 2017 (Core17) [2], and the TREC Washington Post Corpus [^4] https://trec.nist.gov/data/wapost/ used as part of TREC Common Core 2018 (Core18) [34]. Here it is important to note that Core18 includes $\sim$26k ($\sim$4k positive) relevance labels, while Core17 includes $\sim$30k ($\sim$9k positive) labels. For generating queries with Doc2Query, we use the available pre-trained model without finetuning  [^5] https://github.com/terrierteam/pyterrier_doc2query. All of the experiments are run on a Dell workstation with an Intel i9-12900K CPU, 64GB of RAM, and an NVIDIA GeForce RTX 3070 GPU on Ubuntu 22.04 LTS.
@@ 5 Discussion
The results in Section 4.1 show that although the dense retrieval model gives slightly better results, these differences stay the same throughout the simulation. Minor differences in trends align with other studies [13, 33], and in our perspective, user behavior leads to a much greater variance in effectiveness. In particular, taking into account the context within the search session — modeled by dynamic click behavior, background knowledge, and the integration of feedback — substantially impacts information gain.
@@
From our point of view, it is crucial for future simulation experiments to integrate the user context instead of only investigating cost-intensive retrieval models. For both test collections, the trends were similar in all evaluations, even though the absolute information gain differed. We suspect this is related to the substantial difference in the number of documents rated as relevant (cf. 3.5).
@@
In future evaluations, the costs for specific user actions used for the effort-based measure should be specified and justified. Still, the order and trends of the effort-based evaluations resemble the sDCG evaluations closely. Furthermore, the effort-based measure can reveal effects that would remain hidden if only the number of queries were considered. For example, Figure 3 shows that the difference between 50tnr and 20rpp is more pronounced for the effort-based measure since the larger number of examined documents comes with high costs for the user that sDCG does not consider. Regarding sRBP, sDCG also mainly supports the produced orders of the different configurations. However, the information gain stagnates more quickly, and a difference in the orders can be seen in Figure 3, so dynamic click behavior is penalized by sRBP. Both findings can be explained by the substantially stronger discounting of late results by sRBP. Interestingly, dynamic stop behavior is notably rewarded by the other measures. As Figure 4 indicates, the larger number of examined documents can likely further be attributed to this opposing evaluation behavior. Determining not only realistic durations for user actions but also estimating suitable parameters for sRBP is vital for future work to enable more accurate evaluations and could be done by analyzing real user behavior.
@@
The use of probabilistic query generation methods shows promising results. Not only do they dominate rule-based approaches, but they are also more capable of integrating background knowledge of the topic. Integrating feedback into probabilistic approaches would be an exciting direction for future work, as our rule-based approach has clearly shown that feedback is beneficial. Limitations of this work are, on the one hand, a need for estimates of how similar the simulated behavior (e.g., query generation) corresponds to real users. On the other hand, the results obtained cannot be generalized arbitrarily since we only examined a limited variation of simulated users for a specific search task, and both data sets correspond to the same domain.
6 Conclusion
In this work, we propose novel ways to simulate context-dependent query formulations that are evaluated with a state-of-the-art experimental setup, including LLMs and dense retrieval methods. To better understand the effects of context information, we compare context-sensitive query simulation methods against context-free variants. Our experimental results suggest that both the inclusion of comprehensive descriptions of the information need as well as the feedback inferred from earlier seen search results impact the progress of a simulation session, leading to better search effectiveness. In this regard, we envision experiments with user simulations of higher fidelity by considering contextual information.
Acknowledgements This work was supported by Klaus Tschira Stiftung (JoIE - 00.003.2020) and Deutsche Forschungsgemeinschaft (RESIRE - 509543643).